# Regularity and temperature of stationary black hole event horizons

### v1

@@ Abstract
Available proofs of the regularity of stationary black hole event horizons rely on some assumptions on the existence of sections that imply a $C^{1}$ differentiability assumption. By using a quotient bundle approach, we remedy this problem by proving directly that, indeed, under the null energy condition event horizons of stationary black holes are totally geodesic null hypersurfaces as regular as the metric. Only later, by using this result, we show that the cross-sections, whose existence was postulated in previous works, do indeed exist. These results hold true under weak causality conditions. Subsequently, we prove that under the dominant energy condition stationary black hole event horizons indeed admit constant surface gravity, a result that does not require any non-degeneracy assumption, requirements on existence of cross-sections or a priori smoothness conditions. We are able to make sense of the angular velocity and of the value (not just sign) of surface gravity as quantities related to the horizon, without the need of assuming Einstein’s vacuum equations and the Killing extension. Physically, the result means that under very general conditions every stationary black hole has indeed a constant temperature (the zeroth law of black hole thermodynamics).
@@ 1 Stationary black holes
An important problem in mathematical relativity is that of establishing that the event horizon in stationary black holes is as smooth as the metric. Subsequently one can try to show further properties, i.e. that surface gravity is constant, or that the null vector field tangent to the event horizon extends as a Killing vector field (not necessarily coincident with that implied by the initial symmetry) in a neighborhood of the horizon (in the analytic case this is Hawking’s rigidity theorem [16, 1]).
@@
Ideally, such a proof should not rely on any assumption on the smoothness of the event horizon nor on special topological conditions. Horizons are differentiable in the interior of the lightlike geodesic generators and at the endpoints of generators with multiplicity one, but otherwise can be quite non-differentiable [10, 4]. The $C^{1}$ differentiability condition implies that no generator escapes the horizon, that is, every point stays in the interior of a generator. Some further fine properties can be established [8, 9], for instance by using the theory of lower-$C^{2}$ functions [34]. In any case, it has been observed that imposing strong differentiability properties, and possibly even analyticity, on the spacetime manifold, metric, or Cauchy hypersurfaces does not guarantee that horizons will be differentiable. Indeed, an example by Budzyński et al. [6] shows that non-differentiable compact Cauchy horizons may still form.
@@
However, if the null energy condition is imposed then compact Cauchy horizons are indeed as smooth as the metric [34, 30]. Thus, similar behavior can be expected for event horizons, that is, they might likely be shown to be as smooth as the metric under the null energy condition.
@@
Note that the difficulty of the problem lies in the fact that the horizon is not defined locally, e.g. via properties of the Killing field, but rather set theoretically, as the boundary of the past of future infinity. Naturally, it is hard to put one’s hand over such evanishing objects, so technicalities met in works such as [34, 30] devoted to the compact case should be expected. Fortunately, the problem can largely be reduced to an application of the compact case, and also the existence of a Killing symmetry is of help. Strategies that reduce the study of the event horizon for a stationary black holes to that for compact horizons are, of course, not new. The idea was introduced in [16], see also [38], (under smoothness assumptions on the horizon) but, as we shall see, our quotient approach will be different, the quotient being with respect to the isometric flow, not with respect to the geodesic flow of the horizon.
@@
For what concerns available results, a first proof of the smoothness of the event horizon was given in [8, cf. Sec. Conclusions] but details on how to construct certain spacelike sections pushed to the future by the Killing flow were not provided.
@@
More details were given in Chruściel and Costa [7, Thm. 4.1,4.11], where the authors made use of a certain assumption on the existence of cross-sections to lightlike geodesic reaching $\mathscr{I}^{+}$ but, unfortunately, as we shall show, it implies a $C^{1}$ differentiability assumption on the future of such section.
@@
Actually, a certain unsatisfaction in regards to the imposition of the existence of such cross-sections is also found in the original paper by Chruściel and Costa [7, p. 197] where the authors write
@@
We end this section by introducing some definitions and terminology which are the same of [34, 35]. A spacetime $(M,g)$ is a paracompact, time oriented Lorentzian manifold of dimension $n+1\geq 2$. The signature of the metric is (,+,,+)(-,+$\ldots$,+)( - , + , … , + ). We assume that $M$ is $C^{k}$, $4\leq k\leq\infty$, or even analytic, and so as it is $C^{1}$ it has a unique $C^{\infty}$ compatible structure (Whitney) [24, Theor. 2.9]. Thus we could assume that $M$ is smooth without loss of generality.
The metric will be assumed to be $C^{3}$ but it is likely that the degree can be lowered. We assume at least this degree of differentiability because it was also assumed in [34] over which results we rely. A Killing field satisfies $k^{a}{}_{;b;c}=R^{a}{}_{bcd}k^{d}$ and so it is as regular as the metric, and similarly is its flow. For shortness, sometimes we shall sloppily use the word smooth as meaning as much as the regularity of the metric allows when such a regularity is clear.
@@ 2 Existence of the principal bundles
Since $\phi_{t}$ is an homeomorphism $\partial\phi_{t}(S)=\phi_{t}(\partial S)$ (and $\overline{\phi_{t}(S)}=\phi_{t}(\overline{S})$) thus $\partial S=\phi_{t}(\partial S)$ (resp. $\overline{S}=\phi_{t}(\overline{S})$). ∎
We denote with $\phi(S)$ the orbit of the set $S$ as in [16], $\phi(S):=\cup_{t}\phi_{t}(S)$.
@@
Proof.
The last statement follow from the invariance of the Killing flow of $I^{+}(M_{end})$. ∎
Definition 2.3.
@@
Since $H$ is an achronal boundary it has no edge. As a consequence $H^{+}$ might have edge but $\textrm{edge}(H^{+})\cap H^{+}=\emptyset$ (for results on the connection between $\textrm{edge}(H^{+})$, bifurcate horizons and non-degeneracy, see [28, 44, 5]).
@@
Proof.
Consider $\partial I^{-}(\gamma_{p},T)$ in the spacetime $(T,g)$ then $\partial I^{-}(\gamma_{p},T)$ is achronal. Since it is invaraint under the flow of $k$ which is timelike on $T$ then $\partial I^{-}(\gamma_{p},T)=\emptyset$ and so $I^{-}(\gamma_{p},T)=T(p)$. $M_{end}\subset T$ is connected and contains $p$ so it is contained in the connected component of $T$ that contains $p$, namely $T(p)$. ∎
Although there is some freedom in defining $M_{end}$, the resulting set $I^{-}(M_{end})$ is largely independent of it as the next results show. The next result is an improvement of [11, Lemma 3.1].
Lemma 2.5.
@@
Proof.
Let $r\in U\cap M_{end}$, then from lemma 2.4, $M_{end}\subset I^{-}(\gamma_{r})$. Since $U$ is invariant and a past set, then $I^{-}(\gamma_{r})\subset U$. Hence $M_{end}\subset U$ and finally $I^{-}(M_{end})\subset U$ since $U$ is a past set. ∎
The next result is essentially [11, Lemma 3.1] but with weaker assumptions.
Lemma 2.6.
@@
Proof.
It follows from the previous result. ∎
Corollary 2.7.
@@
Proof.
We know that $k$ is timelike on $M_{end}$ and $\Sigma_{end}$ is acausal. Suppose there is a future imprisoned orbit starting from $\Sigma_{end}$, then it accumulates on an imprisoned inextendible causal curve that accumulates on itself [33, 35] (it is not necessarily closed). Such limit curve would still be an integral curve of $k$ hence timelike, which means, by the openness of the chronology relation, that any point of the limit curve belongs to the chronology violating set. As the latter is open, the original orbit intersects it, and by isometry that would give that $\Sigma_{end}$ intersects the chronology violating set too, contradicting its acausality. ∎
We recall that a future $C^{0}$ null hypersurface $H$ is a locally achronal topological embedded hypersurface such that for every $p\in H$ there is a future inextendible lightlike geodesic (called generator) contained in $H$ with past endpoint $p$ [18]. Future $C^{0}$ null hypersurface will also be called past horizons.
Proposition 2.9.
@@
Let $r\in M_{end}$, so that $H=\partial I^{-}(\gamma_{r})$, where $\gamma_{r}$, $\gamma_{r}(0)=r$, is the orbit passing through $r$. Let $p\in H=\partial I^{-}(\gamma_{r})$ then we can find $p_{n}\in I^{-}(\gamma_{r})$, $q_{n}\in\gamma_{r}$, such that $p_{n}\ll q_{n}$, $p_{n}\to p$. We can always redefine $q_{n}$ so that $q_{n}=\gamma_{r}(t_{n})$ with $t_{n}\to\infty$ and $q_{n}$ escaping every compact set. By the limit curve theorem, there is a a future inextendible continuous causal curve $\gamma$ in $\overline{I^{-}(M_{end})}$ starting from $p$. But this continuous causal curve cannot intersect $I^{-}(M_{end})$ otherwise $p\in I^{-}(M_{end})$, which proves that the image of $\gamma$ is contained in $H$. As $H$ is achronal, $\gamma$ is an achronal lightlike geodesic. ∎
@@
A set with the property of the next proposition is called cross-section in [7, Sec. 4.1]. It is used in that paper to prove smoothness or analyticity of the event horizon. It has become a standard assumption in the literature on black holes, see e.g. [29]. Unfortunately, it is so strong that it is essentially equivalent to demanding $C^{1}$ differentiability of $H$ from the outset, as it basically imposes that there are no non-differentiability points.
Proposition 2.10.
If $H$ admits a compact subset $K$ intersected precisely once by every generator, then $H$ is $C^{1}$ on $H\backslash\overline{J^{-}(K)}$.
Proof.
From every non-differentiability point $p$ of a $C^{0}$ future null hypersurface $H$ depart at least two distinct generators [4]. Since every generator intersects $K$ it must be $p\in J^{-}(K)$, thus $H$ is differentiable on the open set $H\backslash\overline{J^{-}(K)}$. For a horizon differentiability on an open set is equivalent to $C^{1}$ regularity [4]. ∎
@@
The next result is a slight improvement over [16, Lemma 2.1]. We note that, save for the property of acausality for $\Sigma_{end}$, we did not impose causality conditions so far.
Proposition 2.11.
Let $(M,g)$ be past distinguishing at $H^{+}$, then $k$ does not vanish on $H^{+}$. Moreover, no point $q\in H$ where past distinction holds is such that $I^{-}(q)\supset I^{-}(M_{end})$.
Proof.
If $k$ vanishes at $q\in H\cap I^{+}(M_{end})$, then $\phi_{t}(q)=q$ for every $t$, hence $E:=\{q\}$ is invariant under the flow. Observe that $I^{-}(E)\cap M_{end}\neq\emptyset$ thus, by Lemma 2.6, we have $I^{-}(q)\supset I^{-}(M_{end})$. This inclusion implies $H\subset\overline{I^{-}(q)}$, hence the points on the future generator starting from $q$ share the same chronological past with $q$. This means violation of past distinction at $q$. ∎
Corollary 2.12.
@@
For $p\in M_{end}$ no point $q\in H$ where past distinction holds is such that $I^{-}(q)\supset\gamma_{p}$.
Proof.
@@
Indeed, as $I^{-}(\gamma_{p})=I^{-}(M_{end})$ we would have otherwise $I^{-}(q)\supset I^{-}(M_{end})$, which contradicts the previous proposition. ∎
Proposition 2.13.
@@
Let $p\in M_{end}$. For each compact subset $K\subset H$ at which strong causality holds we can find $p^{\prime}\in\gamma_{p}$ such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$.
(possibly the assumption can be improved to past distinction)
Proof.
@@
If not the family of closed subsets of the compact set $K$ given by $\{\overline{I^{+}(r)}\cap K:r\in\gamma_{p}\}$ satisfies the finite intersection property, which implies that there is $q$ common to all elements of the family. But then any $q^{\prime}\gg q$ is such that $I^{-}(q^{\prime})\supset\gamma_{p}$ and hence $I^{-}(q^{\prime})\supset I^{-}(M_{end})$. Let $r\neq q$ be a point in the generator starting from $q$, to the future of $q$. Let $x\ll q$, then $r\gg x$, but $r\in\overline{I^{-}(q^{\prime})}$ thus by the openness of $I^{+}$ we can, with a future directed timelike curve, start from $x$ reach a point arbitrarily close to $r$ and finally reach $q^{\prime}$. As $x$ and $q^{\prime}$ can be chosen arbitrarily close to $q$, strong causality is violated at $q$, a contradiction. ∎
@@
Proposition 2.14.
Let $(M,g)$ be strongly causal at $H$. There is an invariant open neighborhood $V$ of $H$ such that, for every $q\in M_{end}$ and for every compact set $K\subset V$ there is $q^{\prime}\in\gamma_{q}$ such that $K\cap\overline{I^{+}(q^{\prime})}=\emptyset$.
In the statement we can replace $H$ with $H^{+}$, it is sufficient to apply the same replacement in the proof.
Proof.
Suppose that for every $p\in M_{end}$ there is $V(p)$ with the property that for every compact set $K\subset V(p)$ there is $p^{\prime}\in\gamma_{p}$ such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$.
@@
Let us prove the assumption in the first paragraph of the proof. Let $p\in M_{end}$ and let $q\in H$. Accordingly to Prop. 2.13 there is $p^{\prime}(q)\in\gamma_{p}$ such that $q\notin\overline{I^{+}(p^{\prime})}$. As this last set is closed we can find an open set $O(q)$ such that $O(q)\cap\overline{I^{+}(p^{\prime})}=\emptyset$. Let us define the invariant open set
$V(p)=\cup\{\phi_{t}(O(q)),q\in H,t\in\mathbb{R}\}=\phi(\cup_{q}O(q)),$
and let $K\subset V(p)$ be compact. There is a finite covering $\{U_{i}\}$ of $K$, $U_{i}:=\phi_{t_{i}}(O(q_{i}))$, to whose elements $O(q_{i})$ correspond points $p^{\prime}_{i}\in\gamma_{p}$ such that $O(q_{i})\cap\overline{I^{+}(p_{i}^{\prime})}=\emptyset$. The last point $p^{\prime}$ of the finite family $\{\phi_{t_{i}}(p^{\prime}_{i})\}$ is then such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$. ∎
Theorem 2.15.

### v2

@@ Abstract
Available proofs of the regularity of stationary black hole event horizons rely on certain assumptions on the existence of sections that imply a $C^{1}$ differentiability assumption. By using a quotient bundle approach, we remedy this problem by proving directly that, indeed, under the null energy condition event horizons of stationary black holes are totally geodesic null hypersurfaces as regular as the metric. Only later, by using this result, we show that the cross-sections, whose existence was postulated in previous works, indeed exist. These results hold true under weak causality conditions. Subsequently, we prove that under the dominant energy condition stationary black hole event horizons indeed admit constant surface gravity, a result that does not require any non-degeneracy assumption, requirements on existence of cross-sections or a priori smoothness conditions. We are able to make sense of the angular velocity and of the value (not just sign) of surface gravity as quantities related to the horizon, without the need of assuming Einstein’s vacuum equations and the Killing extension. Physically, this implies that under very general conditions every stationary black hole has indeed a constant temperature (the zeroth law of black hole thermodynamics).
@@ 1 Stationary black holes
An important problem in mathematical relativity is that of establishing that the event horizon in stationary black holes is as regular as the metric. Subsequently one can try to show further properties, i.e. that surface gravity is constant, or that the null vector field tangent to the event horizon extends as a Killing vector field (not necessarily coincident with that implied by the initial symmetry) in a neighborhood of the horizon (in the analytic case this is Hawking’s rigidity theorem friedrich99 ; alexakis10 ).
@@
Ideally, such a proof should not rely on any assumption on the smoothness of the event horizon nor on special topological conditions. Horizons are differentiable in the interior of the lightlike geodesic generators and at the endpoints of generators with multiplicity one, but otherwise can be quite non-differentiable chrusciel98 ; beem98 . The $C^{1}$ differentiability condition implies that no generator escapes the horizon, that is, every point stays in the interior of a generator (beem98, , Cor. 3.7). Some further fine properties can be established chrusciel01 ; chrusciel02 , for instance by using the theory of lower-$C^{2}$ functions minguzzi14d . In any case, it has been observed that imposing strong differentiability properties, and possibly even analyticity, on the spacetime manifold, metric, or Cauchy hypersurfaces does not guarantee that horizons will be differentiable. Indeed, an example by Budzyński et al. budzynski03 shows that non-differentiable compact Cauchy horizons may still form.
@@
However, if the null energy condition is imposed then compact Cauchy horizons are indeed as regular as the metric minguzzi14d ; larsson14 . Thus, similar behavior can be expected for event horizons, that is, they might likely be shown to be as regular as the metric under the null energy condition.
@@
Note that the difficulty of the problem lies in the fact that the horizon is not defined locally, e.g. via properties of the Killing field, but rather set theoretically, as the boundary of the past of future infinity. Naturally, it is hard to put one’s hand over such evanishing objects, so technicalities met in works such as minguzzi14d ; larsson14 devoted to the compact case should be expected. Fortunately, the problem can largely be reduced to an application of the compact case, and also the existence of a Killing symmetry is of help. Techniques for reducing the analysis of non-compact stationary black hole event horizons to the compact case are not novel. The idea was introduced in friedrich99 , see also moncrief08 , (under smoothness assumptions on the horizon) but, as we shall see, our quotient approach will be different, the quotient being with respect to the isometric flow, not with respect to the geodesic flow of the horizon.
@@
For what concerns available results, a first proof of the smoothness of the event horizon was given in (chrusciel01, , cf. Sec. Conclusions) but details on how to construct certain spacelike sections pushed to the future by the Killing flow were not provided.
@@
More details were given in Chruściel and Costa (chrusciel08, , Thm. 4.1,4.11), where the authors made use of a certain assumption on the existence of cross-sections to lightlike geodesic reaching $\mathscr{I}^{+}$ but, unfortunately, as we shall show, it implies a $C^{1}$ differentiability assumption on the future of such section.
@@
Actually, a certain unsatisfaction in regards to the imposition of the existence of such cross-sections is also found in the original paper by Chruściel and Costa (chrusciel08, , p. 197) where the authors write
@@
The paper is structured as follows. In Section 2, starting from minimal assumptions, we show the existence of a neighborhood $V$ of the event horizon $H^{+}$ with the structure of a trivial $\mathbb{R}$-bundle, the horizon having projection $S$.
In Section 3, by considering the $\mathbb{Z}$-action, we show that $H^{+}$ can be compactified into a horizon $\hat{H}$, provided $S$ is compact. Under the null convergence condition, this allows us to establish the smoothness of the horizon by using previous results for compact horizons.
In Section 4, we establish the existence of cross-sections, that is, compact codimension one submanifolds of the horizon intersected precisely once by every generator.
In Section 5, we consider the quotient of the horizon with respect to the geodesic flow. This identifies a neighborhood of the event horizon with another trivial $\mathbb{R}$-bundle but with a different quotient. In particular, the projection of the event horizon is now a manifold $\tilde{S}$. In spacetime dimension 4 and in other cases of interest, $S$ and $\tilde{S}$ are shown to be homeomorphic.
In Section 6, we provide a definition of sign of surface gravity sufficiently general to apply to the non-compact horizon case and relate it to conditions of completeness for a lightlike field $n$ tangent to the horizon. The field $n$ turns out to be unique up to rescalings. When the scale is fixed, typically via other properties, the surface gravity is defined.
Finally, in Section 7, we establish that under very general conditions, every event horizon admits a constant surface gravity and an angular velocity. Our findings are then compared with previous results in the literature, particularly those relating to the rigidity and black hole uniqueness theorems.
We end this section by introducing some definitions and terminology which are the same of minguzzi14d ; minguzzi18b . A spacetime $(M,g)$ is a paracompact, time oriented Lorentzian manifold of dimension $n+1\geq 2$. The signature of the metric is (,+,,+)(-$\ldots,+$). We assume that $M$ is $C^{k}$, $4\leq k\leq\infty$, or even analytic, and so as it is $C^{1}$ it has a unique $C^{\infty}$ compatible structure (Whitney) (hirsch76, , Theor. 2.9). Thus we could assume that $M$ is smooth without loss of generality.
The metric will be assumed to be $C^{3}$ but it is likely that the degree can be lowered. We assume at least this degree of differentiability because it was also assumed in minguzzi14d over which results we rely. A Killing field satisfies $k^{a}{}_{;b;c}=R^{a}{}_{bcd}k^{d}$ and so it is as regular as the metric, and similarly is its flow. For shortness, sometimes we shall sloppily use the word smooth as meaning as much as the regularity of the metric allows when such a regularity is clear.
With a curve symbol we might denote the map or the image of the map. The inclusion $\subset$ is reflexive. We refer the reader to minguzzi18b for results in causality theory and conventions not explicitly recalled in the present work.
Note added in proof: Our new work (arXiv:2506.20004v2) establishes that the null energy condition suffices wherever the dominant energy condition is used in this paper.
@@ 2 Existence of the principal bundles
Since $\phi_{t}$ is a homeomorphism $\partial\phi_{t}(S)=\phi_{t}(\partial S)$ (and $\overline{\phi_{t}(S)}=\phi_{t}(\overline{S})$) thus $\partial S=\phi_{t}(\partial S)$ (resp. $\overline{S}=\phi_{t}(\overline{S})$). $\square$
We denote with $\phi(S)$ the orbit of the set $S$ as in friedrich99 , $\phi(S):=\cup_{t}\phi_{t}(S)$.
@@
Proof
It follows from Prop. 1. $\square$
Definition 1
@@
Since $H$ is an achronal boundary it has no edge. As a consequence $H^{+}$ might have edge but $\textrm{edge}(H^{+})\cap H^{+}=\emptyset$ (for results on the connection between $\textrm{edge}(H^{+})$, bifurcate horizons and non-degeneracy, see kay91 ; racz92b ; boyer69 ).
@@
Proof
Consider $\partial I^{-}(\gamma_{p},T)$ in the spacetime $(T,g)$ then $\partial I^{-}(\gamma_{p},T)$ is achronal. Since it is invariant under the flow of $k$ which is timelike on $T$ we have $\partial I^{-}(\gamma_{p},T)$ $\emptyset$ and so $I^{-}(\gamma_{p},T)=T(p)$. $M_{end}\subset T$ is connected and contains $p$ so it is contained in the connected component of $T$ that contains $p$, namely $T(p)$. $\square$
Although there is some freedom in defining $M_{end}$, the resulting set $I^{-}(M_{end})$ is largely independent of it as the next results show. The next result is an improvement of (chrusciel94c, , Lemma 3.1).
Lemma 2
@@
Proof
Let $r\in U\cap M_{end}$, then from lemma 1, $M_{end}\subset I^{-}(\gamma_{r})$. Since $U$ is invariant and a past set, then $I^{-}(\gamma_{r})\subset U$. Hence $M_{end}\subset U$ and finally $I^{-}(M_{end})\subset U$ since $U$ is a past set. $\square$
The next result is essentially (chrusciel94c, , Lemma 3.1) but with weaker assumptions.
Lemma 3
@@
Proof
It follows from the previous result. $\square$
Corollary 1
@@
Proof
We know that $k$ is timelike on $M_{end}$ and $\Sigma_{end}$ is acausal. Suppose there is a future imprisoned orbit starting from $\Sigma_{end}$, then it accumulates on an imprisoned inextendible causal curve that accumulates on itself minguzzi07f ; minguzzi18b (it is not necessarily closed). Such limit curve would still be an integral curve of $k$ hence timelike, which means, by the openness of the chronology relation, that any point of the limit curve belongs to the chronology violating set. As the latter is open, the original orbit intersects it, and by isometry that would give that $\Sigma_{end}$ intersects the chronology violating set too, contradicting its acausality. $\square$
We recall that a future $C^{0}$ null hypersurface $H$ is a locally achronal topological embedded hypersurface such that for every $p\in H$ there is a (possibly non-unique) future inextendible lightlike geodesic (called generator) contained in $H$ with past endpoint $p$ galloway00 . Future $C^{0}$ null hypersurface will also be called past horizons.
Proposition 4
@@
Let $r\in M_{end}$, so that $H=\partial I^{-}(\gamma_{r})$, where $\gamma_{r}$, $\gamma_{r}(0)=r$, is the orbit passing through $r$. Let $p\in H=\partial I^{-}(\gamma_{r})$ then we can find $p_{n}\in I^{-}(\gamma_{r})$, $q_{n}\in\gamma_{r}$, such that $p_{n}\ll q_{n}$, $p_{n}\to p$. We can always redefine $q_{n}$ so that $q_{n}=\gamma_{r}(t_{n})$ with $t_{n}\to\infty$ and $q_{n}$ escaping every compact set. By the limit curve theorem, there is a a future inextendible continuous causal curve $\gamma$ in $\overline{I^{-}(M_{end})}$ starting from $p$. But this continuous causal curve cannot intersect $I^{-}(M_{end})$ otherwise $p\in I^{-}(M_{end})$, which proves that the image of $\gamma$ is contained in $H$. As $H$ is achronal, $\gamma$ is an achronal lightlike geodesic. $\square$
Observe that $H^{+}$ is an open invariant subset of $H$. By continuity the Killing orbits cannot leave a connected component $C$ of $H^{+}$ to enter another one, thus the Killing flow sends each component of $H^{+}$ to itself, $\phi(C)=C$.
@@
A set with the property of the next proposition is called cross-section in (chrusciel08, , Sec. 4.1). It is used in that paper to prove smoothness or analyticity of the event horizon. It has become a standard assumption in the literature on black holes, see e.g. kunduri13 . Unfortunately, it is so strong that it is essentially equivalent to demanding $C^{1}$ differentiability of $H$ from the outset, as it basically imposes that there are no non-differentiability points.
Proposition 5
[…]
Proof
[…]
@@
The next result is a slight improvement over (friedrich99, , Lemma 2.1). We note that, save for the property of acausality for $\Sigma_{end}$, we did not impose causality conditions so far.
We recall that past distinction holds at a point $q$ if for all $p\in M$, $I^{-}(p)=I^{-}(q)\Rightarrow p=q$. It also admits an equivalent formulation in terms of the existence of past distinguishing neighborhoods cf. (hawking73, , Sec. 6.4) (minguzzi06c, , Rem. 3.12).
Proposition 6
From every points $q\in\overline{I^{-}(M_{end})}$ starts a future inextendible continuous causal curve $\sigma$ entirely contained in $\overline{I^{-}(M_{end})}$ (if $q\in H$, by achronality, this is necessarily a generator of $H$ passing from $q$).
@@
No point $q\in\overline{I^{-}(M_{end})}$ where past distinction holds is such that $I^{-}(q)\supset I^{-}(M_{end})$.
If past distinction holds at $q\in\overline{I^{-}(M_{end})}\cap I^{+}(M_{end})$ then $k$ does not vanish at $q$.
Proof
Let $r\in M_{end}$. We know that $I^{-}(\gamma_{r})=I^{-}(M_{end})$ and from Prop. 3 $\gamma_{r}$ escapes every compact set in the future. As $q\in\overline{I^{-}(M_{end})}$ we can find $t_{k}\to+\infty$ and $q_{k}\to q$ such that $(q_{k},\gamma_{r}(t_{k}))\in I$. By the limit curve theorem there is a future inextendible continuous causal curve $\sigma$ starting from $q$ and entirely contained in $\overline{I^{-}(M_{end})}$.
Suppose $q$ as in the second statement exists and let $p\in\sigma$, $p\neq q$, be in the curve $\sigma\subset\overline{I^{-}(M_{end})}$ constructed in the first paragraph. Then $I^{-}(q)\subset I^{-}(p)$ and, as $p$ belongs to the closure of the past set $I^{-}(M_{end})$, $I^{-}(p)\subset I^{-}(M_{end})\subset I^{-}(q)$, a contradiction with past distinction at $q$.
For the second statement, if $k$ vanishes at $q\in\overline{I^{-}(M_{end})}\cap I^{+}(M_{end})$, then $\phi_{t}(q)=q$ for every $t$, hence $E:=\{q\}$ is invariant under the flow. Observe that $I^{-}(E)\cap M_{end}\neq\emptyset$ thus, by Lemma 3, we have $I^{-}(q)\supset I^{-}(M_{end})$, which, due to the previously proved result, gives a contradiction. $\square$
Corollary 2
For $p\in M_{end}$ no point $q\in\overline{I^{-}(M_{end})}$ where past distinction holds is such that $I^{-}(q)\supset\gamma_{p}$.
Proof
@@
Indeed, as $I^{-}(\gamma_{p})=I^{-}(M_{end})$ we would have otherwise $I^{-}(q)\supset I^{-}(M_{end})$, which contradicts the previous proposition. $\square$
Proposition 7
@@
Let $p\in M_{end}$. For each compact subset $K\subset\overline{I^{-}(M_{end})}$ at which strong causality holds we can find $p^{\prime}\in\gamma_{p}$ such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$.
Proof
@@
If not the family of closed subsets of the compact set $K$ given by $\{\overline{I^{+}(r)}\cap K:r\in\gamma_{p}\}$ satisfies the finite intersection property, which implies that there is $q$ common to all elements of the family. But then any $q^{\prime}\gg q$ is such that $I^{-}(q^{\prime})\supset\gamma_{p}$ and hence $I^{-}(q^{\prime})\supset I^{-}(M_{end})$.
As shown in Prop. 6, there is a future inextendible continuous causal curve $\sigma$ starting from $q$ and entirely contained in $\overline{I^{-}(M_{end})}$. Let $r\neq q$ be a point in $\sigma$, to the future of $q$. Let $x\ll q$, then $r\gg x$, but $r\in\overline{I^{-}(q^{\prime})}$ thus by the openness of $I^{+}$ we can, with a future directed timelike curve, start from $x$ reach a point arbitrarily close to $r$ and finally reach $q^{\prime}$. As $x$ and $q^{\prime}$ can be chosen arbitrarily close to $q$, strong causality is violated at $q$, a contradiction. $\square$
@@
Proposition 8
Let $(M,g)$ be strongly causal on an invariant set $A\subset\overline{I^{-}(M_{end})}$. There is an invariant open set $V\supset A$ such that, for every $q\in M_{end}$ and for every compact set $K\subset V$ there is $q^{\prime}\in\gamma_{q}$ such that $K\cap\overline{I^{+}(q^{\prime})}=\emptyset$.
This result can be applied with $A$ replaced by $H$, $H^{+}$, any connected component of $H^{+}$, or $\overline{I^{-}(M_{end})}$ itself. In the following propositions we shall only consider the $H^{+}$ case, though generalizations are possible.
Proof
Suppose that for every $p\in M_{end}$ there is $V(p)$, invariant open set containing $A$, with the property that for every compact set $K\subset V(p)$ there is $p^{\prime}\in\gamma_{p}$ such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$.
@@
Let us prove the assumption in the first paragraph of the proof. Let $p\in M_{end}$ and let $q\in A$. Accordingly to Prop. 7 there is $p^{\prime}(q)\in\gamma_{p}$ such that $q\notin\overline{I^{+}(p^{\prime})}$. As this last set is closed we can find an open set $O(q)$ such that $O(q)\cap\overline{I^{+}(p^{\prime})}=\emptyset$. Let us define the invariant open set
$V(p)=\cup\{\phi_{t}(O(q)),q\in A,t\in\mathbb{R}\}=\phi(\cup_{q}O(q)),$
and let $K\subset V(p)$ be compact. There is a finite covering $\{U_{i}\}$ of $K$, $U_{i}:=\phi_{t_{i}}(O(q_{i}))$, to whose elements $O(q_{i})$ correspond points $p^{\prime}_{i}\in\gamma_{p}$ such that $O(q_{i})\cap\overline{I^{+}(p_{i}^{\prime})}=\emptyset$. The last point $p^{\prime}$ of the finite family $\{\phi_{t_{i}}(p^{\prime}_{i})\}$ is then such that $K\cap\overline{I^{+}(p^{\prime})}=\emptyset$. $\square$
Theorem 2.1